OpenStudy (anonymous):

1. Is the set of integers closed under division? 2. Is the set of perfect squares closed under multiplication? 3. Is the set of odd integers closed under addition?

OpenStudy (anonymous):

1. No, 1 / 7 is not an integer. 2. Yes: x^2 * y^2 = (xy)^2 3. No, 3 + 3 = 6, which is even.